# Characterizations of graph classes via convex geometries: A survey

### v6

@@ 4 Other convexities
In this section we present some contributions to the study of convex geometries. Subsections 4.3 and 4.4 provide examples of convex geometries not associated with path systems.
@@ 4.1 $P_{3}$ convexity
Consider the path system that maps a graph $G$ to its paths of length two (not necessarily induced). The convexity defined over such paths is called the $P_{3}$ convexity of $G$ campos-et-al , and the associated convex sets are called $P_{3}$ convex sets. Note that $S$ is a $P_{3}$ convex set if and only if every $x\notin S$ has at most one neighbor in $S$. Thus, the determination of the convex hull of a set in the $P_{3}$ convexity may be viewed as an infection process where noninfected vertices having two or more infected neighbors become infected (for more details, see campos-et-al and references therein).
@@
First, suppose that the $P_{3}$ convexity of $G$ is a convex geometry. If there are vertices $a,b,c\in V(G)$ inducing a $K_{3}$, then $a\in H(\{b,c\})$ and $b\in H(\{a,c\})$, which contradicts the anti-exchange property, since $\{c\}$ is a convex set. Therefore, $G$ is $K_{3}$-free and the endpoints of any edge of $G$ form a convex set. We also have that $G$ is $C_{4}$-free, because if there is an induced $C_{4}=abcda$ in $G$ then $S=\{c,d\}$ is a convex set, $a\in H(S\cup\{b\})$, and $b\in H(S\cup\{a\})$, which contradicts the anti-exchange property. Now, suppose that $G$ contains an induced $C_{5}=abcdea$ in $G$. Since $\{c,d\}$ is a convex set and $G$ is $C_{4}$-free, we have $H(\{c,e\})=\{c,d,e\}$, because if there is $v\in H(\{c,e\})\setminus\{c,d,e\}$ then $\{c,d,e,v\}$ would be an induced $C_{4}$. Since $a\in H(S\cup\{b\})$ and $b\in H(S\cup\{a\})$, the anti-exchange property does not hold. Therefore, $G$ is $C_{5}$-free. Next, suppose that $G$ contains an induced $P_{4}=abcd$. We have that $S=\{a,d\}$ is a convex set, because if there is a $P_{3}=aed$ in $G$ then we would have that $be,ce\not\in E(G)$ (since $G$ is $K_{3}$-free) and that $abcde$ is an induced $C_{5}$, which is not possible. But then the anti-exchange property does not hold because $b\in H(\{a\}\cup S)$ and $c\in H(\{b\}\cup S)$. Thus, $G$ is $P_{4}$-free.
@@ 4.2 Triangle path convexity
We say that a path $P$ in graph $G$ is a triangle path if every two vertices of $P$ with distance greater than 2 in $P$ are non-adjacent in $G$. A set $S\subseteq V(G)$ is tp-convex if for any $x,y\in S$, every vertex belonging to some triangle path between $x$ and $y$ belongs to $S$. The convexity associated with the tp-convex sets of $G$ is the triangle path convexity of $G$ changat ; dourado-sampaio .
@@
First, suppose that the triangle path convexity of $G$ is a convex geometry. If there are vertices $a,b,c\in V(G)$ inducing a $K_{3}$ then $a\in H(\{b,c\})$ and $b\in H(\{a,c\})$, which contradicts the anti-exchange property, since $\{c\}$ is a convex set. Therefore, $G$ is $K_{3}$-free and the endpoints of any edge of $G$ form a convex set. We also have that $G$ is $C_{k}$-free for $k\geq 4$, because if there is an induced $C_{k}=v_{1}\ldots v_{k}v_{1}$ in $G$ ($k\geq 4$) then $S=\{v_{2},v_{3}\}$ is a convex set, $v_{1}\in H(S\cup\{v_{4}\})$, and $v_{4}\in H(S\cup\{v_{1}\})$, which contradicts the anti-exchange property. Hence, $G$ is acyclic.
@@ 4.3 $\mathcal{F}$-free convexities
Let $H$ be a nontrivial graph. Given a graph $G$, we say that $S\subseteq V(G)$ is $H$-free convex if for every $S^{\prime}\subseteq S$ with $|S^{\prime}|=|V(H)|-1$ and $x\in V(G)$, if the subgraph of $G$ induced by $S^{\prime}\cup\{x\}$ is isomorphic to $H$ then $x\in S$. Given a family ${\cal F}$ of nontrivial graphs, we say that $S\subseteq V(G)$ is ${\cal F}$-free convex if $S$ is $H$-free convex for every $H\in{\cal F}$. The convexity associated with the ${\cal F}$-free convex sets of $G$ is the ${\cal F}$-free convexity of $G$.
@@
Let ${\cal F}$ be the family of odd cycles. Then the ${\cal F}$-free convexity of $G$ is a convex geometry if and only if $G$ is bipartite.
@@
Let ${\cal F}$ be the family of graphs that can be obtained from $K_{5}$ or $K_{3,3}$ by subdivision of edges. Then the ${\cal F}$-free convexity of $G$ is a convex geometry if and only if $G$ is planar.
4.4 $P_{4}^{+}$-convexity
Recall that a cograph is a $P_{4}$-free graph. Given a graph $G$, we say that $S\subseteq V(G)$ is $P^{+}_{4}$-convex if, for every induced $P_{4}=abcd$ in $G$, if $a,b,d\in S$ then $c\in S$. The convexity associated with the $P^{+}_{4}$-convex sets of $G$ is the $P_{4}^{+}$-convexity of $G$.
@@ 4.4 $P_{4}^{+}$-convexity
Proof.
Follows from Theorem 23. ∎
@@ 5 Conclusions
In this survey we presented characterizations of several classes of graphs via convex geometries. The characterizations follow the following structure: A graph $G$ is in class ${\mathscr{G}}$ if and only if the ${\mathscr{C}}$-convexity of $G$ is a convex geometry. The table below summarizes the results.
@@
weakly polarizable $m^{3}$
@@
forests triangle path
@@
An interesting question is to increase the table below by including other important classes, such as distance-hereditary graphs and split graphs. Another line of research is to investigate convex geometries defined on digraphs.

### v7

@@ 4 Other convexities
In this section, we present some contributions to the study of convex geometries, some of which not associated with path systems.
@@ 4.1 $P_{3}$ convexity
Consider the path system that maps a graph $G$ to its paths of length two (not necessarily induced). The convexity defined over such paths is called the $P_{3}$ convexity of $G$ [11], and the associated convex sets are called $P_{3}$ convex sets. Note that $S$ is a $P_{3}$ convex set if and only if every $x\notin S$ has at most one neighbor in $S$. Thus, the determination of the convex hull of a set in the $P_{3}$ convexity may be viewed as an infection process where noninfected vertices having two or more infected neighbors become infected (for more details, see [11] and references therein).
@@
First, suppose that the $P_{3}$ convexity of $G$ is a convex geometry. If there are vertices $a,b,c\in V(G)$ inducing a $K_{3}$, then $a\in H(\{b,c\})$ and $b\in H(\{a,c\})$, which contradicts the anti-exchange property, since $\{c\}$ is a convex set. Therefore, $G$ is $K_{3}$-free and the endpoints of any edge of $G$ form a convex set. We also have that $G$ is $C_{4}$-free, because if there is an induced $C_{4}=abcda$ in $G$, then $S=\{c,d\}$ is a convex set, $a\in H(S\cup\{b\})$, and $b\in H(S\cup\{a\})$, which contradicts the anti-exchange property. Now, suppose that $G$ contains an induced $C_{5}=abcdea$ in $G$. Since $\{c,d\}$ is a convex set and $G$ is $C_{4}$-free, we have $H(\{c,e\})=\{c,d,e\}$, because if there is $v\in H(\{c,e\})\setminus\{c,d,e\}$ , then $\{c,d,e,v\}$ would be an induced $C_{4}$. Since $a\in H(S\cup\{b\})$ and $b\in H(S\cup\{a\})$, the anti-exchange property does not hold. Therefore, $G$ is $C_{5}$-free. Next, suppose that $G$ contains an induced $P_{4}=abcd$. We have that $S=\{a,d\}$ is a convex set, because if there is a $P_{3}=aed$ in $G$, then we would have that $be,ce\not\in E(G)$ (since $G$ is $K_{3}$-free) and that $abcde$ is an induced $C_{5}$, which is not possible. But then the anti-exchange property does not hold because $b\in H(\{a\}\cup S)$ and $c\in H(\{b\}\cup S)$. Thus, $G$ is $P_{4}$-free.
@@ 4.2 Triangle path convexity
We say that a path $P$ in graph $G$ is a triangle path if every two vertices of $P$ with distance greater than 2 in $P$ are non-adjacent in $G$. A set $S\subseteq V(G)$ is tp-convex if for any $x,y\in S$, every vertex belonging to some triangle path between $x$ and $y$ belongs to $S$. The convexity associated with the tp-convex sets of $G$ is the triangle path convexity of $G$ [12, 18].
@@
First, suppose that the triangle path convexity of $G$ is a convex geometry. If there are vertices $a,b,c\in V(G)$ inducing a $K_{3}$, then $a\in H(\{b,c\})$ and $b\in H(\{a,c\})$, which contradicts the anti-exchange property, since $\{c\}$ is a convex set. Therefore, $G$ is $K_{3}$-free and the endpoints of any edge of $G$ form a convex set. We also have that $G$ is $C_{k}$-free for $k\geq 4$, because if there is an induced $C_{k}=v_{1}\ldots v_{k}v_{1}$ in $G$ ($k\geq 4$), then $S=\{v_{2},v_{3}\}$ is a convex set, $v_{1}\in H(S\cup\{v_{4}\})$, and $v_{4}\in H(S\cup\{v_{1}\})$, which contradicts the anti-exchange property. Hence, $G$ is acyclic.
@@
4.3 All-path convexity
A set $S\subseteq V(G)$ is ap-convex if, for any $x,y\in S$, every vertex lying in a path between $x$ and $y$ belongs to $S$. The convexity associated with the ap-convex sets of $G$ is the all-path convexity of $G$. A study on the all-path convexity can be found in [39].
It is easy to see that $v$ is an extreme vertex of an ap-convex set $S$ with $|S|>1$ if and only if $v$ is a pendant vertex (a vertex of degree one) in $G[S]$.
Theorem 27.
The all-path convexity of a graph $G$ is a convex geometry if and only if $G$ is a forest.
Proof.
Suppose that the all-path convexity of $G$ is a convex geometry and $G$ contains a cycle $C$. Then, the vertices in $V(C)$ are non-pendant and $H(V(C))$ contains no extreme vertices, a contradiction. Thus $G$ is a forest. Conversely, if $G$ is a forest, then every non-pendant vertex of $G$ lies in a path between two pendant vertices, and this property holds for every induced subgraph of $G$. This means that every ap-convex set of $G$ is the convex hull of its extreme vertices. ∎
4.4 Cycle convexity
We say that a set $S\subseteq V(G)$ is c-convex if no vertex $u\notin S$ is adjacent to distinct $x,y\in S$ such that there is a path from $x$ to $y$ in $G[S]$; in other words, no $u\notin S$ is contained in a cycle $C$ with $V(C)\backslash S=\{u\}$. The convexity associated with the c-convex sets of $G$ is the cycle convexity of $G$. See [4, 5] for important results on the cycle convexity.
Clearly, $x$ is an extreme vertex of a c-convex set $S$ if and only if there is no cycle $C$ in $G[S]$ containing $x$.
Theorem 28.
The cycle convexity of a graph $G$ is a convex geometry if and only if $G$ is a forest.
Proof.
[…]
It is interesting to note that Theorems 26,  27, and 28 all apply to the class of forests. However, the triangle path, all-path, and cycle convexities do not coincide in general. If $G=C_{n}$ ($n\geq 4$), the families of tp-convex sets, ap-convex sets, and c-convex sets of $G$ are distinct.
4.5 $\Delta$-convexity
A set $S\subseteq V(G)$ is $\Delta$-convex if no vertex $u\notin S$ has two adjacent neighbors in $S$, that is, forms a triangle with two vertices in $S$. The $\Delta$-convexity of $G$ is the family formed by the $\Delta$-convex sets of $G$ (see [3] for more details). It is easy to see that $x$ is an extreme vertex of a $\Delta$-convex set $S$ if and only if $N(x)$ is an independent set of $G[S]$.
Theorem 29.
The $\Delta$-convexity of a graph $G$ is a convex geometry if and only if $G$ is triangle-free.
Proof.
[…]
@@ 4.6 $\mathcal{F}$-free convexities
We now introduce the concept of $\mathcal{F}$-free convexity [7]. Let $H$ be a nontrivial graph. Given a graph $G$, we say that $S\subseteq V(G)$ is $H$-free convex if for every $S^{\prime}\subseteq S$ with $|S^{\prime}|=|V(H)|-1$ and $x\in V(G)$, if the subgraph of $G$ induced by $S^{\prime}\cup\{x\}$ is isomorphic to $H$, then $x\in S$. Given a family ${\cal F}$ of nontrivial graphs, we say that $S\subseteq V(G)$ is ${\cal F}$-free convex if $S$ is $H$-free convex for every $H\in{\cal F}$. The convexity associated with the ${\cal F}$-free convex sets of $G$ is the ${\cal F}$-free convexity of $G$.
@@
Let ${\cal F}$ be the family of odd cycles. Then, the ${\cal F}$-free convexity of $G$ is a convex geometry if and only if $G$ is bipartite.
@@
Let ${\cal F}$ be the family of graphs that can be obtained from $K_{5}$ or $K_{3,3}$ by subdivision of edges. Then, the ${\cal F}$-free convexity of $G$ is a convex geometry if and only if $G$ is planar.
The ${\cal F}$-free convexity defined by =${\cal F}=\{P_{4}\}$ is called $P_{4}^{+}$-convexity. Recall that a cograph is a $P_{4}$-free graph. Given a graph $G$, we say that $S\subseteq V(G)$ is $P^{+}_{4}$-convex if, for every induced $P_{4}=abcd$ in $G$, if $a,b,d\in S$, then $c\in S$. The convexity associated with the $P^{+}_{4}$-convex sets of $G$ is the $P_{4}^{+}$-convexity of $G$.
@@
To conclude this section, we remark that the cycle convexity and the $\Delta$-convexity are ${\cal F}$-free convexities defined by taking =${\cal F}=\{C_{k}\mid k\geq 3\}$ and =${\cal F}=\{C_{3}\}$, respectively. Hence, Theorems 28 and 29 can be viewed as corollaries of Theorem 30.
@@ 5 Conclusions
In this survey, we presented characterizations of several classes of graphs via convex geometries. The characterizations follow the following structure: A graph $G$ is in class ${\mathscr{G}}$ if and only if the ${\mathscr{C}}$-convexity of $G$ is a convex geometry. The table below summarizes the results.
An interesting question is to increase the table by including other important classes, such as distance-hereditary graphs and split graphs. Another line of research is to investigate convex geometries defined on digraphs. Finally, as far as the authors now, the $g^{3}$-Steiner convexities that are convex geometries have not been characterized yet.
@@
weak bipolarizable $m^{3}$
Ptolemaic Steiner
$(P_{4},G_{1},G_{2},G_{3},G_{4})$-free 3-Steiner
(Theorem 10) ${\cal R}$-Steiner
(Theorem 11) $g^{3}_{3}$-Steiner
@@
forest triangle path, all path, cycle
triangle-free $\Delta$